OpenStudy (anonymous):

Simplify the expression. State any restrictions on the variable. x^2+6x+8/x^2-10x-24

OpenStudy (anonymous):

\[x^2+6x+8 \over x^2-10x-24\]

OpenStudy (inkyvoyd):

Factor the top andbottom. Tell me what you get.

OpenStudy (anonymous):

x^2-10x-24 = (x-12)(x+2)

OpenStudy (anonymous):

x^2+6x+8 = (x+2)(x+4)